# Multi-Agent Trajectory Prediction with Difficulty-Guided Feature Enhancement Network

### v1

@@ III-F Future feature enhancement
$\hat{F}=\mathbf{MLP}(\mathbf{flatten}(\hat{P}_{\text{masked}}))$ (17)
Subsequently, we interact the future trajectory feature $\hat{F}$ with the agent-centric interactive actor feature $\hat{A}^{(k)}$ via multi-head attention. This operation aims to impose certain constraints on the prediction of other agents through the reliable future trajectory features. Finally, we need to perform one last interaction using the original map features $\hat{M}$. This step is crucial as it aims to refocus on the reachable lane lines of the map after imposing social constraints through the future trajectory features[34]. These two operations can be formulated as:
$H=\mathbf{MHA}_{AF}(\hat{A}^{(k)},\hat{F})$ (18)
$\hat{O}=\mathbf{MHA}_{AM}(H,\hat{M})$ (19)
@@ III-G Final prediction
${A}_{fusion}=\text{Cat}(\hat{O},\hat{X}^{(l)})$ (20)
$E=\mathbf{MLP}_{\text{end}}({A}_{fusion})+\mathbf{MLP}_{\text{refine}}(\text{Cat}[{A}_{fusion},E])$ (21)
$K=\mathbf{softmax}(\mathbf{MLP}_{\text{cls}}(\text{Cat}[A_{fusion},E]))$ (22)
$P=\text{Cat}[\mathbf{MLP}_{\text{traj}}(\text{cat}[{A}_{fusion},E]),E]$ (23)
Here, $P$ represents the multimodal prediction trajectory and its corresponding probability $K$ that we finally obtain.
@@
Model Training. Although our feature enhancement component outputs trajectories during the intermediate process, our training process remains end-to-end. Our DGFNet obtains initial and final predictions in a single pass. Similar to the previous method [29, 22, 12, 25], we supervise the output trajectories $P$ and $\hat{P}$ during the training process by smoothing the L1 loss, which is expressed as ${\mathcal{L}}_{reg}$ and ${\mathcal{L}}_{reg}^{c}$. For the categorization loss in $K$, we use the maximum marginal loss, which is expressed as ${\mathcal{L}}_{cls}$. Our overall loss expression is as follows:
@@ IV-A Experimental setup
Datasets. Our model DGFNet is tested and evaluated on a very challenging and widely used self-driving motion prediction dataset: the Argoverse 1&2 [47, 48]. Both motion prediction datasets provide agent tracking trajectories and semantically rich map information at a frequency of 10Hz over a specified time interval. The prediction task in Argoverse 1 is to predict trajectories for the next 3 seconds based on the previous 2 seconds of historical data. The dataset contains a total of 324,557 vehicle trajectories of interest extracted from over 1000 hours of driving. In contrast, Argoverse 2 requires a higher level of robustness and generalization of the predictive model, with 250,000 of the most challenging scenarios officially filtered from the self-driving test fleet. Argoverse 2 predicts the next 6 seconds from the first 5 seconds of historical trajectory data, and Argoverse 2 provides heading data for traffic targets as well as traffic target categories, in order to ensure fair comparisons between models, the dataset has official data partitioning and the test set is evaluated using the Eval online test server.
Evaluation Metrics. We have adopted the standard testing and evaluation methodology used in motion prediction competitions to assess prediction performance. Key metrics for individual agents include Probabilistic minimum Final Displacement Error (p-minFDE), Minimum Final Displacement Error (minFDE), Minimum Average Displacement Error (minADE), Miss Rate (MR) and Drivable Area Compliance (DAC). Where p-minFDE, MR, and minFDE reflect the accuracy of the predicted endpoints, and minADE indicates the overall bias in the predicted trajectories. DAC, reflects the compliance of the predicted outcomes, which is incorporated into the bias when the predicted outcomes are outside of the drivable region.
Implementation Details. We generate lane vectors for lanes that are more than 50 meters away from any available agent. The number of layers $l$ and $k$ of interaction features are set to 3. All layers except the final decoder have 128 output feature channels. In addition, the weight parameters in the loss function are set to $\alpha$=0.7, $\beta$=0.1 and $\lambda$=0.2. The hyperparameter $\tau$ in the difficulty masker was set to 5. The model was trained 50 epochs on an Nvidia RTX 3090 with a batch size of 16. For the experiments on Argoverse 1&2, we used a segmented constant learning rate strategy: up to the 5th epoch we used $5\times 10^{-5}$; from the 6th epoch to the 25th, we used $5\times 10^{-4}$; from the 26th epoch to the 40th epoch, we used $3\times 10^{-4}$; and thereafter, we used $5\times 10^{-5}$ until training was complete. Similar to LaneGCN[29], data augmentation was achieved by applying random scaling in the range [0.75, 1.25].
@@ IV-B Quantitative Results
Comparison with State-of-the-Arts in Agoverse 1. We conducted performance comparisons between our method and the state-of-the-art methods on the test set of Argoverse 1 over the past three years. It should be noted that in the table, LaneGCN [29], DenseTNT [35], THOMAS [36], SceneTransformer [33], HiVT-128 [25], LAformer [14], GANet [14], R-Pred [49] performed inference as single models. HOME+GOHOME [41, 42], Multipath++ [43], DCMS [44], Wayformer [45], ProphNet [16], QCNet [46], and FFINet [17] used ensemble inference involving 5 or 10 models. Where Macformer [38] and HeteroGCN [40] have both single and ensemble inference results.
According to the supplementary materials of relevant papers, the ensemble methods used for the results with model ensembles are inconsistent. Generally, ensemble methods involve using K-means to match the closest trajectories from multiple models, followed by weighted computation to obtain the final result[50]. We obtained 5 models by setting different initial random seeds and performed ensemble inference using the aforementioned method. The results are shown in Tab.I. Compared to single-model methods, our approach consistently ranks at the top in almost all metrics. Even when compared with methods using ensemble strategies, our approach remains highly competitive, achieving prediction accuracy comparable to the leading models such as QCNet and ProphNet in the leaderboard. However, our method has advantages in terms of parameter size and inference time, which will be detailed in the following sections.
@@
Comparison with State-of-the-Arts in Agoverse 2. In our comparison of performance with several recent single-model inference methods on the Argoverse 2 test set, we found that while the advantage of our method is not as pronounced as in Argoverse 1, it remains reasonably acceptable. We speculate that this is primarily due to the increased prediction horizon in the Argoverse 2 dataset, which has led to less robust initial prediction results. QCNet[46] was the first to propose mitigating the instability of long-term predictions through recurrent step. In future work, we will also explore similar approaches to improve robustness. Moreover, due to the inclusion of categories such as pedestrians and cyclists in Argoverse 2, the effectiveness of the parameter settings for our difficulty masker has also been impacted. Overall, our model demonstrates highly competitive performance within an acceptable range of model parameters across both datasets.
Comparison of Prediction Performance on High-difficulty Samples. Tab.III shows the highest minFDE errors of each method on the Argoverse 1 validation set. We reasoned on the validation set with LaneGCN’s pre-trained model to obtain some sample sets with the highest errors, and then reasoned on these sample sets using our base model and DGFNet, respectively. The results show that our method DGFNet has significantly lower minFDE errors than the comparison methods in all percentage ranges, which fully demonstrates the significant advantage of our method in mitigating the long-tail problem in trajectory prediction.
Quantitative Results and Visualization. Tab.4 shows the qualitative visualization of DGFNet on the Argoverse 1&2 validation set. The model provides reasonable multimodal predictions for multiple target agents in different traffic scenarios. The eight selected scenarios have high prediction difficulty and strong representativeness. In Seq.1, Seq.2, and Seq.3, the model generates accurate predictions for each agent in a busy and strongly interacting intersection scenario. In Seq.3 and Seq.5, the model gives robust predictions in the face of vehicle turning tendencies. As a comparison, in Seq.6, Seq.7 and Seq.8, the model gives more scattered prediction results when the intention of the vehicle is not clear. In addition, the model captures the final landing very accurately in both turning and straight ahead situations.
TABLE IV: The ablation study results of different components on the Argoverse 1 validation split, including Difficulty Masker (DM), Future Feature Interaction (FFI), and Scene Feature Fusion (SFF).
@@ IV-C Analysis and Discussion
Ablation Study. We conducted an ablation study on the Argoverse 1 validation set, which consists of 39,472 sequences. The initial model only used agent-centric historical trajectories and maps for feature extraction. The second model incorporated scene-centric features into the initial model, achieving some performance improvement through Scene Feature Fusion (SFF). The third model added the Future Feature Interaction (FFI) module, enabling interaction between extracted historical features and future features. The final model used the Difficulty Masker (DM) to mask out future trajectories that are difficult to predict, ensuring that only reliable trajectories are used for future interaction. Of concern is that without adding Difficulty Masker, the performance of direct future enhancement drops dramatically, failing because the second stage model may find shortcuts during optimization if the results given in the first stage are good enough [51]. It tends to replicate the results of the first stage rather than trying to extract richer information from future features. According to Tab.IV, comparing the base model with DGFNet, the improvement in predictive performance is evident. And we also did ablation experiments on the effect of parameter $\tau$ on the model. According to Tab.V, the model performance obtained by setting parameter $\tau$ to 5 is the best.
@@
Computational Performance. We compared the parameters of each method and the average inference time per scenario on the same device. As shown in Tab.VI, our inference time is slightly higher than LaneGCN and close to the lighter models HiVT-128 and LAformer. However, our method has a much higher prediction accuracy than these two methods. Compared with the methods with the highest prediction accuracy, our method has fewer parameters (note that those marked with * use ensemble strategies), which means that our model achieves high prediction accuracy with lower computational cost. Then, we recorded the inference time on the same experimental machine equipped with an NVIDIA GeForce RTX 3090. Specifically, we adjusted the batch size of all models to 1 and recorded their average time on the validation set. The model inference time indicates that the real-time latency of DGFNet is within an acceptable range. Since LAformer uses a two-stage inference approach, its inference time is longer despite having fewer parameters than our method.
@@ V Conclusion
In this paper, we propose a novel Difficulty-Guided Feature Enhancement Network (DGFNet) for muti-agent trajectory prediction. Distinguishing from general future enhancement networks, our model emphasizes filtering future trajectories by masking out and retaining only reliable future trajectories for feature enhancement, which greatly improves the prediction performance. Extensive experiments on the Argoverse 1&2 benchmarks show that our method outperforms most state-of-the-art methods in terms of prediction accuracy and real-time processing. In addition, we integrate agent-centric and scence-centric scene features in a concise and effective way, and demonstrate that the fusion of the two features is effective in improving the model prediction accuracy.
Emulating the predictive habits of human drivers is an intriguing direction for future research. This involves emulating human strategies when encountering different vehicles and making effective assumptions in the presence of incomplete perceptual information. Furthermore, we will integrate Difficulty-Guided feature enhancement into other prediction frameworks to further validate the efficacy of this module in optimizing multi-agent trajectory prediction problems.
Figure 4: Quantitative results of DGFNet on the Argoverse 1&2 validation set: observed trajectories of the proxy past are highlighted in orange, future trajectories of the ground truth are highlighted in red, and predicted trajectories are shown in green. The top four scenarios correspond to the Argoverse 1 dataset and the bottom four scenarios correspond to the Argoverse 2 dataset.

### v2

@@ III-F Future feature enhancement
^$\hat{\mathcal{F}}=\mathbf{MLP}(\mathbf{flatten}(\hat{\mathcal{P}}_{\text{masked}}))$ (17)
Subsequently, we interact the future trajectory feature $\hat{F}$ with the agent-centric interactive actor feature $\hat{\mathcal{A}}^{(k)}$ via multi-head attention. This operation aims to impose certain constraints on the prediction of other agents through the reliable future trajectory features. Finally, we need to perform one last interaction using the original map features $\hat{\mathcal{M}}$. This step is crucial as it aims to refocus on the reachable lane lines of the map after imposing social constraints through the future trajectory features[23]. These two operations can be formulated as:
$\mathcal{H}=\mathbf{MHA}_{AF}(\hat{\mathcal{A}}^{(k)},\hat{\mathcal{F}})$ (18)
$\hat{\mathcal{O}}=\mathbf{MHA}_{AM}(\mathcal{H},\hat{M})$ (19)
@@ III-G Final prediction
${\mathcal{A}}_{fusion}=\text{cat}(\hat{\mathcal{O}},\hat{\mathcal{X}}^{(l)})$ (20)
$\mathcal{E}=\mathbf{MLP}_{\text{end}}({\mathcal{A}}_{fusion})+\mathbf{MLP}_{\text{refine}}(\text{cat}[{\mathcal{A}}_{fusion},\mathcal{E}])$ (21)
$\mathcal{K}=\mathbf{softmax}(\mathbf{MLP}_{\text{cls}}(\text{cat}[\mathcal{A}_{fusion},\mathcal{E}]))$ (22)
$\mathcal{P}=\text{cat}[\mathbf{MLP}_{\text{traj}}(\text{cat}[{\mathcal{A}}_{fusion},\mathcal{E}]),\mathcal{E}]$ (23)
here, $\mathcal{P}$ represents the multimodal prediction trajectory and its corresponding probability $\mathcal{K}$ that we finally obtain.
@@
Model Training. Although our feature enhancement component outputs trajectories during the intermediate process, our training process remains end-to-end. Our DGFNet obtains initial and final predictions in a single pass. Similar to the previous method [16, 17, 1, 13], we supervise the output trajectories $P$ and $\hat{\mathcal{P}}$ during the training process by smoothing the $\mathcal{L}1$ loss, which is expressed as ${\mathcal{L}}_{reg}$ and ${\mathcal{L}}_{reg}^{c}$. For the categorization loss in $\mathcal{K}$, we use the maximum marginal loss, which is expressed as ${\mathcal{L}}_{cls}$. Our overall loss expression is as follows:
@@ IV-A Experimental setup
Datasets. Our model DGFNet is tested and evaluated on a very challenging and widely used self-driving motion prediction dataset: the Argoverse 1&2 [32, 33]. Both motion prediction datasets provide agent tracking trajectories and semantically rich map information at a frequency of 10Hz over a specified time interval. The prediction task in Argoverse 1 is to predict trajectories for the next 3 seconds based on the previous 2 seconds of historical data. The dataset contains a total of 324,557 vehicle trajectories of interest extracted from over 1000 hours of driving. Argoverse 2 contains 250,000 of the most challenging scenarios officially filtered from the self-driving test fleet. Argoverse 2 predicts the next 6 seconds from the first 5 seconds of historical trajectory data. To ensure fair comparisons between models, both datasets were subjected to official data partitioning and the test set was evaluated using the Eval online test server.
Evaluation Metrics. We have adopted the standard testing and evaluation methodology used in motion prediction competitions to assess prediction performance. Key metrics for individual agents include Probabilistic minimum Final Displacement Error (p-minFDE), Minimum Final Displacement Error (minFDE), Minimum Average Displacement Error (minADE), Miss Rate (MR) and Drivable Area Compliance (DAC). Where p-minFDE, MR, and minFDE reflect the accuracy of the predicted endpoints, and minADE indicates the overall bias in the predicted trajectories. DAC, reflects the compliance of the predicted outcomes.
Implementation Details. We generate lane vectors for lanes that are more than 50 meters away from any available agent. The number of layers $l$ and $k$ of interaction features are set to 3. All layers except the final decoder have 128 output feature channels. In addition, the weight parameters in the loss function are set to $\alpha$=0.7, $\beta$=0.1 and $\lambda$=0.2. The hyperparameter $\tau$ in the difficulty masker was set to 5. The model was trained 50 epochs on an Nvidia RTX 3090 with a batch size of 16. For the experiments on Argoverse 1&2, we used a segmented constant learning rate strategy: up to the 5th epoch we used $5\times 10^{-5}$; from the 6th epoch to the 40th, we used $5\times 10^{-4}$; and thereafter, we used $5\times 10^{-5}$ until training was complete.
@@ IV-B Quantitative Results
Comparison with State-of-the-Arts in Agoverse 1. We conducted performance comparisons on the Argoverse 1 test set using both single-model and multi-model approaches between our method and the most classical and latest state-of-the-art methods. Generally, ensemble methods involve using K-means to match the closest trajectories from multiple models, followed by weighted computation to obtain the final result. We obtained 5 models by setting different initial random seeds and performed ensemble inference using the aforementioned method. The results are shown in Tab.I. Compared to single-model methods, our approach consistently ranks at the top in almost all metrics. Even when compared with methods using ensemble strategies, our approach remains highly competitive, achieving prediction accuracy comparable to the leading models such as QCNet and ProphNet in the leaderboard. However, our method has advantages in terms of parameter size and inference time, which will be detailed in the following sections.
@@
Comparison with State-of-the-Arts in Agoverse 2. In our comparison of performance with several recent single-model inference methods on the Argoverse 2 test set, we found that while the advantage of our method is not as pronounced as in Argoverse 1, it remains reasonably acceptable. We speculate that this is primarily due to the increased prediction horizon in the Argoverse 2 dataset, which has led to less robust initial prediction results. Moreover, due to the inclusion of categories such as pedestrians and cyclists in Argoverse 2, the effectiveness of the parameter settings for our difficulty masker has also been impacted. Overall, our model demonstrates highly competitive performance within an acceptable range of model parameters across both datasets.
@@
Comparison of Prediction Performance on High-difficulty Samples. Tab.III shows the highest minFDE errors of each method on the Argoverse 1 validation set. We reasoned on the validation set with LaneGCN’s pre-trained model to obtain some sample sets with the highest errors, and then reasoned on these sample sets using our base model and DGFNet, respectively. It can be observed that the average accuracy difference for all samples is minimal compared to Our base. However, our method, DGFNet, has significantly lower minFDE errors across all percentage ranges, clearly demonstrating the remarkable advantage of our method in mitigating the long-tail problem in trajectory prediction.
Quantitative Results and Visualization. Fig.4 shows the qualitative visualization. The four selected scenarios have high prediction difficulty and strong representativeness. In Seq.1 and Seq.2, the model generates accurate predictions for each agent in a busy and strongly interacting intersection scenario. In Seq.3, the model gives robust predictions in the face of vehicle turning tendencies. As a comparison, in Seq.4 the model gives more scattered prediction results when the intention of the vehicle is not clear. Overall, the model captured the final landing point very accurately in both turning and straight ahead situations.
Table IV: The ablation study results of different components on the Argoverse 1 validation set.
@@ IV-C Analysis and Discussion
Ablation Study. We conducted an ablation study on the Argoverse 1 validation set. The initial model only used agent-centric historical trajectories and maps for feature extraction. The second model incorporated scene-centric features into the initial model, achieving some performance improvement through Scene Feature Fusion (SFF). The third model added the Future Feature Interaction (FFI) module, enabling interaction between extracted historical features and future features. The final model used the Difficulty Masker (DM) to mask out future trajectories that are difficult to predict, ensuring that only reliable trajectories are used for future interaction. Of concern is that without adding Difficulty Masker, the performance of direct future enhancement drops dramatically, failing because the second stage model may find shortcuts during optimization if the results given in the first stage are good enough [34]. According to Tab.IV, comparing the base model with DGFNet, the improvement in predictive performance is evident. And we also did ablation experiments on the effect of parameter $\tau$ on the model. According to Tab.V, the model performance obtained by setting parameter $\tau$ to 5 is the best.
@@
Computational Performance. We compared the parameters of each method and the average inference time per scenario on the same device. As shown in Tab.VI, our inference time is slightly higher than LaneGCN and close to the lighter models HiVT-128 and LAformer. However, our method has a much higher prediction accuracy than these two methods. Compared with the methods with the highest prediction accuracy, our method has fewer parameters (note that those marked with * use ensemble strategies), which means that our model achieves high prediction accuracy with lower computational cost. Then, we recorded the inference time on the same experimental machine equipped with an NVIDIA GeForce RTX 3090. The model inference time indicates that the real-time latency of DGFNet is within an acceptable range. Since LAformer uses a two-stage inference approach, its inference time is longer despite having fewer parameters than our method.
@@ V Conclusion
In this paper, we propose a novel Difficulty-Guided Feature Enhancement Network (DGFNet) for muti-agent trajectory prediction. Distinguishing from general future enhancement networks, our model emphasizes filtering future trajectories by masking out and retaining only reliable future trajectories for feature enhancement, which greatly improves the prediction performance. Extensive experiments on the Argoverse 1&2 benchmarks show that our method outperforms most state-of-the-art methods in terms of prediction accuracy and real-time processing. Emulating the predictive habits of human drivers is an intriguing direction for future research. This involves emulating human strategies when encountering different vehicles and making effective assumptions in the presence of incomplete perceptual information.
Figure 4: Quantitative results of DGFNet on the Argoverse 1&2 validation set. The top scenarios correspond to the Argoverse 1 dataset and the bottom scenarios correspond to the Argoverse 2 dataset.